OpenStudy (anonymous):

(4-5/6+3*2)/(5/6) please with steps thank you very much :)

OpenStudy (anonymous):

Some more brackets may be handy here. Assuming (4 - (5/6) + 3*2)/(5/6): = (19/6 + 6)/(5/6) = (55/6)/(5/6) = 55/5 = 11